# Can LLMs Replace Human Evaluators? An Empirical Study of LLM-as-a-Judge in Software Engineering

### v2

@@ 2.2. SE Benchmarks and Metrics
Many SE benchmarks focus solely on code generation, where LLMs generate code for the given requirements and function signatures. HumanEval (Chen et al., 2021) is one of the most adopted code generation benchmarks, featuring 164 human-curated Python problems. It uses Pass@$k$ as the evaluation metric. MBPP (Austin et al., 2021) is another popular benchmark with 974 Python problems, aiming at entry-level developers. APPS (Hendrycks et al., 2021) is a much larger Python benchmark with 10000 problems, ranging from being solvable in one-line to presenting substantial challenges in algorithms. ClassEval (Du et al., 2023) challenges LLMs with 100 class-level code generation problems in Python, and measure class-level and method-level Pass@$k$.
@@ 3. LLM-as-a-judge Framework Overview
    LLMs can also be fine-tuned as specialized judges, usually applied with a single inference pass and no additional strategies. State-of-the-art LLMs like GPT-4 are often used to generate the reference judgment for training. In this paper, we discuss the performance of these LLMs, instead of focusing on the detailed training process.
@@ 4. Study Design
In this section, we elaborate on the details of our study design. In our study, we focus on leveraging different types of LLM-as-a-judge methods to evaluate the responses of three SE tasks. We collect instructions and generate responses from representative datasets, and then perform human and LLM evaluation on these responses, and analyze their correlations.
@@ 4.1.2. Response Generation
After preliminary experiments, we discover that many reference summaries in CodeXGLUE are in fact incorrect. Therefore, we require the reference summary to at least have 15 tokens, reselect the instructions, and manually examine each instruction. We also find that in code generation, selected LLMs struggle to generate interpretable code because they cannot use dependencies effectively, as the only available dependency information in ComplexCodeEval is their names, which makes human evaluation almost impossible to yield meaningful scores. Consequently, we reselect the instructions with at most 5 dependencies to reduce difficulty, and augment the dependency information with GPT-4o [^8] We use the 2024-08-06 version for all experiments., prompting [^9] The prompt for augmentation is available in our repository. it to extract the signature from the reference answer [^10] The authors of ComplexCodeEval extract dependency names from reference answers as well, and we follow their practice to utilize reference answers. and generate a short description for each dependency. We manually examine the descriptions to ensure that no other information about the reference answers is included. Responses are generated with the updated dependency information as well as other contextual information.
@@ 4.1.3. Manual Evaluation
Two human evaluators with expertise in the chosen programming languages are involved in judging each of the 450 responses. During manual evaluation, we provide the corresponding instruction and the reference answer along with the response to be evaluated. Each evaluator is required to score both aspects before assigning an overall score, which is not necessarily the average of the former. The final human score for each response is the average of overall scores from two evaluators [^11] The two human evaluators reach a high level of agreement, achieving Spearman’s $\rho$ of (83.07, 75.42, 74.20), Pearson’s $R$ of (85.86, 79.70, 73.74), and Kendall’s $\tau$ of (72.26, 63.40, 62.57) on code translation, generation, and summarization respectively.. For pairwise comparison, we calculate the absolute difference between the final human scores of two responses in a pair, declaring a tie when the difference is smaller than 0.5121212The value is chosen so that ties occur for about a third of the response pairs for each task., or deciding the higher-scored response is better otherwise.
@@ 4.2.3. Probability-based Methods
We select two probability-based methods. These methods may take at least two of the following as input: instruction $src$, response $tgt$, and reference $ref$, plus supplementary information like evaluation aspects [^13] Aspects are identical to those in human evaluation.. We use davinci-002 here, since later OpenAI models only return probabilities of newly generated tokens instead of provided tokens.
@@ 4.2.4. Output-based Methods
G-Eval (Liu et al., 2023) requires the LLM to generate the evaluation steps first and embeds it into the prompt, followed by 20 inference passes with a high temperature of 1.0 and averaging the scores. Following their practice, we prompt the LLM to return the score first with a limit of 20 generated tokens. Judgments without a score are discarded. For pairwise comparison, we consider comparison results as scores of 1 or -1 when one response is better, or 0 for a tie. If the absolute value of the average score is less than 0.7151515The value is chosen so that ties occur for about a third of the response pairs., we declare a draw, otherwise considering one response better.
@@ 5.1. RQ1: Alignment with Human Scores
On the contrary, LLM-as-a-judge methods struggle to outperform conventional metrics in evaluating code generation outputs and are completely surpassed in evaluating code summarization. For code generation, conventional metrics can reach a mid-high correlation of $\rho=67.11$, $R=65.55$, and $\tau=49.66$, while DeepSeek-V2.5 is the only LLM outperforming them at $\rho=66.39$, $R=68.51$, and $\tau=54.74$ without any additional inference strategies. This can be attributed to the characteristics of the ComplexCodeEval dataset, which emphasizes the usage of complicated dependencies by filling out the correct arguments and calling them at the right time instead of designing sophisticated algorithms. Therefore, a response-reference comparison at the lexical-level can offer an insight of the response’s quality, while the LLMs’ limited understanding of the dependencies fail to provide benefits in evaluation. With that said, for code generation, LLM-as-a-judge methods with large LLMs like GPT-4o are still applicable, since they display similar performance as conventional metrics but provide the benefits of not requiring reference answers. For code summarization, LLM-as-a-judge techniques are completely defeated by conventional metrics, hardly reaching a score of 30 in any correlation coefficient or even demonstrating a negative correlation with human evaluation. Nonetheless, conventional metrics also fail to deliver satisfying alignment with human evaluation, with $\rho,R<50$ and $\tau<40$. This is potentially due to the fact that many LLMs try to explain the code step-by-step instead of summarizing the core functionality, which is difficult for these LLM-as-a-judge methods to detect. While conventional metrics can assign low scores to these responses, they have trouble handling paraphrasing, which is common in summaries. It is an interesting future direction to explore new metrics that align with humans for code summarization.
@@
Similarly, current inference strategies, when employed to GPT-4o, produce an inadequate performance boost of $\Delta R=2.21,6.03,3.04$ at maximum. Despite recent work claiming the effectiveness of scaling inference (Snell et al., 2024), we found that existing inference strategies for SE evaluation only bring marginal improvement in human alignment. Moreover, they have different downsides: G-Eval forces LLMs to generate the overall score, restricting the efficacy of the Chain-of-Thought procedure, while greatly increasing inference cost if the full explanations are needed; BatchEval increases the token count, leading to more expensive inference due to multi-round evaluation. Therefore, greedy decoding remains a viable LLM-as-a-judge solution with satisfactory performance and lower requirements of token count, when equipped with colossal state-of-the-art LLMs.
@@ 5.2. RQ2: Score Characteristics
Table 5. Correlation between output-based LLM-as-a-judge methods grouped by the sizes of the LLMs they use. Methods are categorized based on the underlying model size: ”Small” (using ¡50B LLMs), which includes all methods from the SFT group and DeepSeek-Coder-V2-Lite from the Vanilla group, and ”Large” (using ¿230B LLMs), including DeepSeek-V2.5 and GPT-4o, the latter used by G-Eval and BatchEval.
@@
Output-based methods using large LLMs tend to align well with each other, whereas those using smaller LLMs exhibit low correlations with other methods. Since output-based methods offer the best human alignment, we further investigate whether LLM size influences correlations between methods by grouping these LLM-as-a-judge methods into those using large LLMs (¿230B) and those using small LLMs (¡50B). In Table 5, we observe that methods employing large LLMs achieve high correlations of $\rho>80$ for code translation and $\rho>65$ for code generation with each other. These methods use DeepSeek-V2.5 and GPT-4o, and maintain strong alignment despite the difference in LLMs and inference strategies. In contrast, methods using small LLMs yield $\rho<50$ when compared to methods in the ”large” group, and $\rho<30$ within the ”small” group. This pattern reflects a performance gap, as the ”large” group align substantially better with human evaluations than the ”small” group.
@@ 5.3. RQ3: Pairwise Comparison versus Individual Scoring
Table 6. Experimental results for pairwise comparison, where Acc. means Accuracy and Agr. means agreement. The best alignment in each column is marked in bold. Results better than random guessing are underlined.

### v3

@@ 2.2. SE Benchmarks and Metrics
Many SE benchmarks focus solely on code generation, where LLMs generate code for the given requirements and function signatures. HumanEval (Chen et al., 2021) is one of the most adopted code generation benchmarks, featuring 164 human-curated Python problems. It uses Pass@$k$ as the evaluation metric. MBPP (Austin et al., 2021) is another popular benchmark with 974 Python problems, aiming at entry-level development. APPS (Hendrycks et al., 2021) is a much larger Python benchmark with 10000 problems, ranging from being solvable in one-line to presenting substantial challenges in algorithms. ClassEval (Du et al., 2023) challenges LLMs with 100 class-level code generation problems in Python, and measure class-level and method-level Pass@$k$.
@@ 3. LLM-as-a-Judge Framework Overview
    LLMs can also be fine-tuned as specialized judges, usually applied with a single inference pass and no additional strategies. State-of-the-art LLMs like GPT-4 are often used to generate reference judgments for training. In this paper, we discuss the performance of these LLMs, instead of focusing on the detailed training process.
@@ 4. Study Design
In this section, we elaborate on the details of our study design. In our study, we focus on leveraging different types of LLM-as-a-judge methods to evaluate the responses of three SE tasks. We collect instructions and generate responses from representative datasets, then perform human and LLM evaluation on these responses, and analyze their correlations.
@@ 4.1.2. Response Generation
After preliminary experiments, we discover that many reference summaries in CodeXGLUE are in fact incorrect. Therefore, we require the reference summary to at least have 15 tokens, reselect the instructions, and manually examine each instruction. We also find that in code generation, selected LLMs struggle to generate interpretable code because they cannot use dependencies effectively, as the only available dependency information in ComplexCodeEval is their names, which makes human evaluation almost impossible to yield meaningful scores. Consequently, we reselect the instructions with at most five dependencies to reduce difficulty, and augment the dependency information with GPT-4o [^8] We use the 2024-08-06 version for all experiments. The prompt for augmentation is available in our repository., prompting it to extract the signature from the reference answer [^9] The authors of ComplexCodeEval extract dependency names from reference answers as well, and we follow their practice to utilize reference answers. and generate a short description for each dependency. We manually examine the descriptions to ensure that no other information about the reference answers is included. Responses are generated with the updated dependency information as well as other contextual information.
@@ 4.1.3. Manual Evaluation
Two human evaluators with expertise in the chosen programming languages are involved in judging each of the 450 responses. During manual evaluation, we provide the corresponding instruction and the reference answer along with the response to be evaluated. Each evaluator is required to score both aspects before assigning an overall score, which is not necessarily the average of the former. The final human score for each response is the average of overall scores from two evaluators [^10] The two human evaluators reach a high level of agreement, achieving Spearman’s $\rho$ of (83.07, 75.42, 74.20), Pearson’s $R$ of (85.86, 79.70, 73.74), and Kendall’s $\tau$ of (72.26, 63.40, 62.57) on code translation, generation, and summarization respectively.. For pairwise comparison, we calculate the absolute difference between the final human scores of two responses in a pair, declaring a tie when the difference is smaller than 0.5111111The value is chosen so that ties occur for about a third of the response pairs for each task., or deciding the higher-scored response is better otherwise.
@@ 4.2.3. Probability-Based Methods
We select two probability-based methods. These methods may take at least two of the following as input: instruction $src$, response $tgt$, and reference $ref$, plus supplementary information like evaluation aspects [^12] Aspects are identical to those in human evaluation. $a$. We use davinci-002 here, since later OpenAI models only return probabilities of newly generated tokens instead of prompt tokens.
@@ 4.2.4. Output-Based Methods
G-Eval (Liu et al., 2023) requires the LLM to generate the evaluation steps first and embeds it into the prompt, followed by 20 inference passes with a high temperature of 1.0 and averaging the scores. Following their practice, we prompt the LLM to return the score first with a limit of 20 generated tokens. Judgments without a score are discarded. For pairwise comparison, we consider comparison results as scores of 1 or -1 when one response is better, or 0 for a tie. If the absolute value of the average score is less than 0.7141414The value is chosen so that ties occur for about a third of the response pairs., we declare a draw, otherwise considering one response better.
@@ 5.1. RQ1: Alignment with Human Scores
On the contrary, LLM-as-a-judge methods struggle to outperform conventional metrics in evaluating code generation outputs and are completely surpassed in evaluating code summarization. For code generation, conventional metrics can reach a mid-high correlation of $\rho=67.11$, $R=65.55$, and $\tau=49.66$, while DeepSeek-V2.5 is the only LLM outperforming them at $\rho=66.39$, $R=68.51$, and $\tau=54.74$ without any additional inference strategies. This can be attributed to the characteristics of the ComplexCodeEval dataset, which emphasizes the usage of complicated dependencies by filling out the correct arguments and calling them at the right time instead of designing sophisticated algorithms. Therefore, a response-reference comparison at the lexical level can offer an insight of the response’s quality, while the LLMs’ limited understanding of the dependencies fail to provide benefits in evaluation. With that said, for code generation, LLM-as-a-judge methods with large LLMs like GPT-4o are still applicable, since they display similar performance as conventional metrics but provide the benefits of not requiring reference answers. For code summarization, LLM-as-a-judge techniques are completely defeated by conventional metrics, hardly reaching a score of 30 in any correlation coefficient or even demonstrating a negative correlation with human evaluation. Nonetheless, conventional metrics also fail to deliver satisfying alignment with human evaluation, with $\rho,R<50$ and $\tau<40$. This is potentially due to the fact that many LLMs try to explain the code step-by-step instead of summarizing the core functionality, which is difficult for these LLM-as-a-judge methods to detect. While conventional metrics can assign low scores to these responses, they have trouble handling paraphrasing, which is common in summaries. It is an interesting future direction to explore new metrics that align with humans for code summarization.
@@
Similarly, current inference strategies, when employed to GPT-4o, produce an inadequate performance boost of $\Delta R=2.21,6.03,3.04$ at maximum. Despite recent work claiming the effectiveness of scaling inference (Snell et al., 2024), we found that existing inference strategies for SE evaluation only bring marginal improvement in human alignment. Moreover, they have different downsides: G-Eval forces LLMs to generate the overall score first, restricting the efficacy of the Chain-of-Thought procedure, while greatly increasing inference cost if the full explanations are needed; BatchEval increases the token count, leading to more expensive inference due to multi-round evaluation. Therefore, greedy decoding remains a viable LLM-as-a-judge solution with satisfactory performance and lower requirements of token count, when equipped with colossal state-of-the-art LLMs.
@@ 5.2. RQ2: Score Characteristics
Table 5. Correlation between output-based LLM-as-a-judge methods grouped by the sizes of the LLMs they use. Methods are categorized based on the underlying model size: ”Small” (using ¡50B LLMs), which includes all methods from the SFT group and DeepSeek-Coder-V2-Lite from the Vanilla group, and ”Large” (using ¿100B LLMs), including DeepSeek-V2.5 and GPT-4o, the latter used by G-Eval and BatchEval.
@@
Output-based methods using large LLMs tend to align well with each other, whereas those using smaller LLMs exhibit low correlations with other methods. Since output-based methods offer the best human alignment, we further investigate whether LLM size influences correlations between methods by grouping these LLM-as-a-judge methods into those using large LLMs (¿100B) and those using small LLMs (¡50B). In Table 5, we observe that methods employing large LLMs achieve high correlations of $\rho>80$ for code translation and $\rho>65$ for code generation with each other. These methods use DeepSeek-V2.5 and GPT-4o, and maintain strong alignment despite the difference in LLMs and inference strategies. In contrast, methods using small LLMs yield $\rho<50$ when compared to methods in the ”large” group, and $\rho<30$ within the ”small” group. This pattern reflects a performance gap, as the ”large” group align substantially better with human evaluations than the ”small” group.
@@ 5.3. RQ3: Pairwise Comparison versus Individual Scoring
Table 6. Experimental results for pairwise comparison, where Acc. means Accuracy and Agr. means agreement. The best result in each column is marked bold. Results higher than 50 are underlined.